# Tint Your Models Task-wise for Improved Multi-task Model Merging

### v1

@@ 4.1 Model Tinting
Loss correlation. Here, we justify our design choice as represented in the loss in Eq. (2). Inspired by the analysis [65], we demonstrate the effectiveness of using cross-entropy with model-generated labels by computing the Spearman correlation between cross-entropy losses based on individual model signals and ground-truths; the correlation of AdaMerging’s entropy loss versus the ground-truth-based cross-entropy loss is also computed. In addition, we further examine the correlation before and after training to see how training affects the correlation.
Table 1 reports the correlation coefficients for each task. We use the weight-averaged model [19] to provide baseline values in the ‘before’ rows, while ‘after’ denotes the correlations measured using a trained merged model. We observe that the entropy loss initially shares the correlations with the supervised cross-entropy loss, and the correlation mostly dropped after training across diverse tasks (i.e., 1 out of 8), particularly on challenging ones like Cars, even showed a negative correlation. This suggests that entropy minimization may not fully follow the true supervisory signals for MTL [^3] As also observed in prior studies [65, 38], entropy sometimes diverges further from the supervised cross-entropy for certain sample distributions.; the resulting merged model does not appear to be trained under ground-truth supervisory signals.
In contrast, our model-generated cross-entropy loss initially shows higher correlations and maintains a striking, strong positive correlation after training. Intriguingly, the correlation strengthens after training, which suggests that the merged model becomes more aligned with supervisory signals, likely due to its improvement under appropriately emulated signals. Therefore, our approach effectively trains a merged model that can better follow the ground-truth signals at test time when labels are inaccessible.
4.2 Model Tinting++
Improving Model Tinting. Model-generated supervisory signals are crucial for our method, and enhancing these signals could yield further improvements. Our insight is to exclude undesirable samples during training where the merged model is correct while individual models are incorrect in that training with such signals may not always be beneficial. How many such samples exist, and what is their impact? Figure 5 shows the extent to which the number of such cases is not negligible. Another question could be: how can we identify these samples for exclusion?
First, we presume only confident samples for individual models would be meaningful in training. Additionally, we refer to the literature [15], which suggests that a calibrated model achieves robustness by mitigating overconfidence; we similarly assume that a merged model would follow the trend and less-confident samples would be helpful [38]. Since quantifying confidence is challenging and introducing a new confidence threshold is undesirable, we sample based on the confidence trends in both individual and merged models. Specifically, we use the top-1 probability as the metric and filter out the instances when the merged model’s top-1 probability for a given prediction is higher than that of the individual model.
This approach can be understood as selecting only the samples that could calibrate the merged model while individual models provide reliable supervisory signals for training. This ensures that the resultant merged model becomes more robust, as similarly discussed in [39]; thereby, we expect improved MTL performance across different tasks. Our experimental results will demonstrate the improvements achieved by the proposed method.
To address this, we experiment with a sequential update strategy, where updates are applied immediately after each task’s forward pass instead of the typical one. Interestingly, this sequential update approach gives a positive byproduct: performance improvements over the traditional approach. While the original AdaMerging achieves an average accuracy of 80.1%, our sequential update strategy improves it to 81.6. Similarly, our method also enjoys a gain when using sequential updates (89.6% compared to 89.4%). We argue that this improvement may result from mitigating catastrophic forgetting in continual learning [26] by training tasks sequentially rather than simultaneously.
@@ 4.2 Model Tinting++
Table 3: Multi-task performance comparison across 8 vision tasks. We compare our methods with the competing methods for merging ViT-B/32 (left) and ViT-L/14 (right) fine-tuned models.
@@ 5.1 Experimental Setup
Datasets and metrics. We follow the standard setups in DARE [66] and EMR-Merging [18] for NLP experiments: employing fine-tuned weights for 8 tasks from the GLUE benchmark [54]. Evaluation on each task is performed with task-specific metrics: Matthews correlation for CoLA [56], Pearson and Spearman correlations for STS-B [4], and accuracy for SST-2 [47], MRPC [11], QQP [20], MNLI [57], QNLI [44], and RTE [13]. For vision tasks, we follow the standard Task Arithmetic [19] and AdaMerging [65] setups, using fine-tuned weights on 8 standard image classification tasks. Furthermore, we extend our vision task experiments with 12 more tasks reaching the overall experimental scope to 20 tasks, following [55]: CIFAR-10 [28], CIFAR-100 [28], FER2013 [14], Flowers102 (Flowers) [37], Oxford-IIIT Pet (Pet) [40], PCAM [53], STL10 [8], EMNIST [9], FashionMNIST (FMNIST) [60], Food101[2], KMNIST [7], and Rendered SST-2 (R-SST) [42]. All tasks use the accuracy metric. Additionally, we provide results for a selected subset of 14 out of the 20 tasks in Appendix.
Models. We employ RoBERTa-base [31] for the NLP tasks in our experiments. We use the publicly released weights [^4] https://github.com/harveyhuang18/EMR_Merging by the prior work [18]. For a fair comparison, we run experiments for TWIN-Merging [32] and our method using fine-tuned weights from EMR-Merging. Our method applies Mean Squared Error (MSE) loss to STS-B instead.
For vision experiments, we utilize the pre-trained CLIP [43] weights [^5] https://github.com/mlfoundations/open_clip, specifically ViT-{B/32, B/16, L/14}. Most analyses and ablation studies are performed with ViT-B/32 for comparison. For a fair comparison across the 8 tasks, we use the fine-tuned weights [^6] https://github.com/mlfoundations/task_vectors publicly released by Task Arithmetic [19]. In experiments with 14 and 20 vision tasks, CLIP ViT-B/32 is fine-tuned based on the configuration used in Task Arithmetic to obtain individual models. Due to the lack of 14 and 20 tasks experiments in literature, we reproduce Task Arithmetic, MagMax [34], Ties Merging [63], Consensus TA [55], AdaMerging [65], Surgery [64], and EMR-Merging [18]. More detailed experimental setups are provided in Appendix.
5.2 Main Results
Merging 8 NLP tasks. As shown in Table 2, our Model Tinting largely improved over the best method on average. While EMR-Merging suffers from a significant performance drop on CoLA, and TWIN-Merging on STS-B, Model Tinting demonstrates consistent performance close to individual task-specific models across all tasks. We argue that this reveals the limitations of methods that only handle shared representations and shows the superiority of our method that effectively manages task-specific information. For further details, please refer to Appendix.
Merging 8 vision tasks. Table 3 and Table 3 demonstrate that our approach surpasses traditional MTL methods and achieves average performance improvements of 0.7%p and 0.2%p on ViT-B/32 and ViT-L/14, respectively, compared to the previous state-of-the-art merging methods. (For ViT-B/16, we observe an improvement of 0.7%p; more details are provided in Appendix) Additionally, our method achieves performance close to the individual models, even surpassing them on some tasks. When applying sample filtering, we observe a 0.5%p improvement for ViT-B/32, while the performance gain for ViT-L/14 is minimal. This is likely because Model Tinting already performs nearly as well as individual models when using ViT-L/14.
Merging 20 vision tasks. In Table 4, we observe that for KMNIST, both task arithmetic and AdaMerging exhibit notably low performance. Furthermore, on datasets like PCAM and EMNIST, AdaMerging occasionally results in lower performance than task arithmetic. In contrast, our approach shows a significant improvement, with a 19.3%p performance increase over AdaMerging and surpassing state-of-the-art methods by 2.3%p. Notably, for several tasks, our method even exceeds the performance of individual models. With sample selection, we see an additional 0.5%p performance boost, highlighting the effectiveness of our approach as the number of tasks scales.
@@ 5.3 Emprical Analyses
Sparsity analysis of merging coefficients. Inspired by the literature [66, 63, 10] in which redundant parameters degrade performance due to task-specific parameters’ conflicts, we analyze the learned merging coefficients to observe this effect in our method. We explore sparsity in the merging coefficients after training them alone and jointly with the classifier. Figure 3 shows both configurations sequentially and demonstrates that the joint training enjoys striking sparsity. We find 37.2% versus 55.9% of the coefficients are concentrated near zero (i.e., smaller than 1e-5) with the deviated average accuracy of 83.4% versus 89.6% for each configuration. This suggests that training a task-specific layer, like the classifier, along with merging coefficients, reduces task interference by pruning unnecessary parameters, leading to improved accuracy.
(a) Training coefficients only
(b) Joint training
Figure 3: Impact of joint training on coefficient sparsity. We visualize the merging coefficients to display the sparsity in weights. Joint training of merging coefficients with a task-specific layer promotes sparsity strikingly, with 55.9% of coefficients near zero compared to 37.2% when training coefficients only.
Can inner layers serve as task-specific adjustments? One might wonder if our observed improvements are unique to training just the classifier. We disprove this speculation by evaluating other layers beyond the classifier. Specifically, we train each of the 12 transformer layers in a merged encoder, which is performed under a similar setup in the pilot study. Figure 4 shows transformer layers generally achieve performance matched to training the classifier (see the details in Appendix). This suggests that training a single internal transformer layer can do similarly to what the classifier does. Furthermore, identifying the optimal layer could lead to additional performance improvements. Meanwhile, when we train either all layers or groups of layers — early layers (from layers 0 to 5-th) or late layers (from layers 6 to 11-th) together — performance drops significantly; we believe that training many layers at once in MTL scenarios diminishes the task-agnostic knowledge inherently stored in a pre-trained model.
@@
Figure 4: Impact of training different task-specific layers. Performance comparison of merged models when training: 1) a single inner layer, 2) late layers, 3) early layers, and 4) only the classifier. Training a single layer matches the performance of training only the classifier, while training multiple layers reduces accuracy.
(a) Prediction discrepancy
(b) Overall difference
Assesing merging methods via prediction discrepancy. Here, we assess diverse merged models with their sample prediction correctness with respect to the individual models that often serve as the performance upper bound in MTL due to higher accuracy. Figure 5 displays two sets of bar graphs. First, the upper bars represent samples correctly predicted by the individual models but incorrectly by the merged model, where a lower one aligns more with the individual models. Our method achieves the smallest discrepancy, significantly outperforming AdaMerging and Surgery. The lower bars in Figure 5 represent samples misclassified by the individual models but correctly classified by the merged model, representing unexpected gains of the merged model. A higher one here benefits from the merged model’s predictions beyond those of the individual model.
Figure 5 illustrates the overall impact by computing the difference between the penalties and gains shown in the upper and lower bars of Figure 5. First, our improved method presented in Section 4.2 successfully enhances this metric, where a lower one is better. We believe our methods effectively balance preserving the strengths of the individual models while capturing unexpected gains, leading to superior overall performance. We observe that while closely related methods like AdaMerging [65] (using entropy minimization without constraints) and Surgery [64] (performs post-AdaMerging to preserve individual model predictions) also enjoy some unexpected gains, the penalties observed in the upper bars outweigh the benefits, which incur net decreases in the overall metric.
@@
Initialization robustness. Our method benefits from the robustness of the merging coefficients’ initialization. To highlight this, we present training curves under different initialization settings and compare them with those of other model merging methods. As shown in Figure 6, when only the classifier is trained, the performance is sensitive to initialization; however, when both the classifier and the coefficients are trained together, the method demonstrates substantial robustness. Figure 7 shows that when training a Surgery model or the classifier only, starting from the AdaMerging initialization yields higher performance than starting from the Task Arithmetic initialization, which suggests the sensitivity to initialization.
@@ 6 Conclusion
In this paper, we have introduced Model Tinting , a novel approach to multi-task model merging that enhances task-specific knowledge through adaptively learning task-specific layers, departing from traditional methods that rely on shared representations. Model Tinting++ further involved a novel confidence-based sampling method, which has proven favorable. Extensive experiments demonstrated state-of-the-art performance across various domains, including vision and NLP, with particularly significant improvements on the challenging 20-task vision benchmark. Our empirical studies further justified that our design choice of incorporating just a single task-wise learned layer is sensible. Finally, we discovered additional insights: 1) inner layers all play similar roles in task-specific adaptations; 2) training task-specific layers naturally promotes sparsity in merging coefficients; 3) our joint training is insensitive to initializations with improved results. We believe our study could be a foundation for future research in model merging.
@@
  • Appendix C: Additional results for NLP and vision tasks;
@@ Appendix A Overview of Model Tinting
We give a brief overview of our proposed method Model Tinting , a test-time approach for merging multiple fine-tuned models by introducing a single trainable layer for each task. Figure A provides an illustrative overview of our method. At the top, the figure illustrates the merging process, where multiple fine-tuned models ($\phi_{1},\phi_{2},\ldots,\phi_{K}$) are combined into a shared representation ($\phi_{\text{MTL}}$) using merging coefficients. During this process, a trainable task-specific layer ($\phi_{\text{tint}}^{l}$) is introduced at a chosen transformer layer $l$, capturing task-specific information from the task-agnostic representations of the merged encoder. We argued that any layer can be picked for task adjustments before merging, and this is further supported by the graph in Figure A. The results demonstrate that applying Model Tinting, regardless of the chosen layer, leads to significant performance improvements over the baseline. This highlights the robustness and effectiveness of task-specific adjustments in enhancing task-agnostic representations, with more details provided in Appendix D.1. We also proposed Model Tinting++ , which applies a strategy to filter noisy predictions from individual models.
@@ B.1 Training Details
Main experiments. We initialize the layer-wise merging coefficients for all layers to 0.3 by default, following the values used in the previous method [65]. However, for the 14 and 20 vision tasks, the coefficients are initially set to 0.1, as using 0.3 incurred significantly lower Task Arithmetic [19] performance for these tasks, presumably due to the increased number of tasks. We use the Adam [25] optimizer with momentum parameters (0.9, 0.999) to update the coefficients and the classifier. For vision tasks, the learning rate is set to 0.01 when training the classifier and 0.001 when training only the merging coefficients. For NLP tasks, a learning rate of 0.0005 is used, regardless of whether the classifier, the merging coefficients, or both are being updated. Consistent with prior works [65, 50, 64], vision tasks are trained with 16,000 test samples per task over 500 iterations, using a batch size of 32. Similarly, NLP tasks are trained with 32,000 samples per task over 1,000 iterations, also with a batch size of 32. For each task, the loss is computed, and the model parameters are updated immediately following the forward pass for each batch. To enable sequential processing, the order of tasks is randomized.
Pilot study. The aforementioned training details are applied almost identically to the experiments in the pilot study presented in our main paper. As stated in the main paper, only the classifier was chosen (as a straightforward and practical option) to train on the given Task Arithmetic-merged weights for the 8 vision tasks using ViT-B/32. A difference is that, instead of using the test dataset, the training dataset is used in a supervised manner to produce the results closer to the upper bound, in which the difference stands out more. We train the models for only one epoch.

### v2

@@ 4.2 Model Tinting++
Table 4: Multi-task performance comparison across 8 vision tasks. We compare our methods with the competing methods for merging ViT-B/32 (left) and ViT-L/14 (right) fine-tuned models. $\dagger$ denotes reproduced results.
@@ 5.1 Experimental Setup
Datasets and metrics. We follow the standard setups in DARE [75] and EMR-Merging [21] for NLP tasks, using fine-tuned weights for eight tasks from the GLUE benchmark [62]. Task-specific metrics are employed: Matthews correlation for CoLA [65], Pearson and Spearman correlations for STS-B [5], and accuracy for other tasks [53, 12, 24, 66, 49, 14]. For vision tasks, we adopt the setups from Task Arithmetic [23] and AdaMerging [74], using fine-tuned weights on eight image classification tasks. We extend our experiments to 12 additional tasks [32, 15, 42, 45, 61, 9, 10, 69, 3, 8, 47], covering a total of 20 tasks as in [63]. All vision tasks are evaluated using accuracy. Results for a subset of 14 tasks are provided in the Appendix C.3. For dense prediction evaluation, we use NYUv2 [52], which contains 1,449 RGB-D indoor scene images. It includes three tasks: 13-class semantic segmentation, depth estimation, and surface normal estimation. We evaluate them with mIoU/pixel accuracy, absolute/relative error, and mean angular error, respectively.
Models. For NLP tasks, we use RoBERTa-base [35], leveraging publicly available fine-tuned weights [^4] https://github.com/harveyhuang18/EMR_Merging from EMR-Merging [21]. We conduct experiments on TWIN-Merging [36] and our method under the same settings, using Mean Squared Error (MSE) loss for STS-B. For vision experiments, we employ pre-trained CLIP [48] models [^5] https://github.com/mlfoundations/open_clip, specifically ViT-B/32, B/16, L/14. Most analyses use ViT-B/32. To ensure comparability across eight tasks, we use publicly available fine-tuned weights [^6] https://github.com/mlfoundations/task_vectors from Task Arithmetic [23]. For 14- and 20-task experiments, we fine-tune CLIP ViT-B/32 following the Task Arithmetic configuration. For dense prediction, we adopt ResNet-50 [18] as the backbone, following FusionBench [57]. The model is initialized with ImageNet pre-trained weights and fine-tuned for each task. Since prior works lack results on 14- and 20-task settings, we reproduce multiple previous methods [38, 72, 64, 63, 74, 73, 21].
@@
Method Segmentation Depth Estimation Normal
[…]
Model Tinting
49.0 73.0 45.7 18.6 26.4
Model Tinting++
50.2 73.4 45.1 18.6 26.1
Table 6: Multi-task performance across three dense prediction tasks. We compare the performance of merging ResNet50 models fine-tuned on three dense prediction tasks in NYUv2. $\dagger$ denotes reproduced results.
5.2 Main Results
Merging 8 NLP tasks. As shown in Table 3, our Model Tinting largely improved over the best method on average. While EMR-Merging suffers from a significant performance drop on CoLA, and TWIN-Merging on STS-B, Model Tinting demonstrates consistent performance close to individual task-specific models across all tasks. We argue that this reveals the limitations of methods that rely on a fixed shared encoder and shows the superiority of our method, which effectively manages task-specific information from the adaptive shared encoder.
Merging 8 vision tasks. Table 4 and Table 4 show that our method outperforms traditional MTL methods, improving ViT-B/32 and ViT-L/14 by 0.7%p and 0.2%p over state-of-the-art methods (ViT-B/16 improves by 0.7%p; see Appendix C.3). Notably, for ViT-B/32, this is achieved with just 0.5M trainable parameters, far fewer than the 7M used in prior state-of-the-art approach. Additionally, our method achieves performance close to the individual models regardless of task difficulty and even surpasses them on certain tasks. For instance, on DTD, a particularly challenging dataset, other merging methods exhibit a noticeable performance gap compared to individual models, whereas our approach achieves a higher accuracy of 80.2% compared to 79.4%. When applying sample filtering, we observe a 0.5%p improvement for ViT-B/32, while the performance gain for ViT-L/14 is minimal. This is likely because Model Tinting already performs nearly as well as individual models when using ViT-L/14.
Merging 20 vision tasks. We explore the impact of scaling model merging by expanding the task set from 8 to 20, making the problem more complex. To ensure a fair comparison, we reproduce the performance of existing methods under the same experimental conditions. In Table 5, we observe that for KMNIST, both task arithmetic and AdaMerging exhibit notably low performance. Furthermore, on datasets like PCAM and EMNIST, AdaMerging occasionally results in lower performance than task arithmetic. In contrast, our approach shows a significant improvement, with a 19.3%p performance increase over AdaMerging and surpassing state-of-the-art methods by 2.3%p. Notably, for several tasks, our method even exceeds the performance of individual models. With sample selection, we see an additional 0.5%p performance boost, highlighting the effectiveness of our approach as the number of tasks scales.
Merging dense prediction tasks. Dense prediction tasks, such as segmentation, depth estimation, and normal estimation, pose significant challenges due to their inter-task heterogeneity. Despite this, most model merging studies focus on classification, leaving these tasks largely unexplored. Table 6 presents results for these tasks, highlighting the limitations of previous methods. Surgery, which performs comparably to individual models on eight classification tasks, suffers substantial performance drops in depth and normal estimation. Similarly, EMR-Merging, despite leveraging task-specific dynamic masks, struggles to maintain segmentation performance. In contrast, our approach effectively balances these tasks, maintaining performance close to individual models even in dense prediction settings.
@@ 5.3 Emprical Analyses
Effect of adjusting different layers. A natural question is whether the improvements observed in Model Tinting are specific to training the classifier. To explore this, we analyze the effect of training different layers beyond the classifier by updating merging coefficients along with each of the 12 transformer layers in the merged encoder. As shown in Figure 4, training a single internal layer achieves accuracy comparable to training the classifier, suggesting that task-specific information can be effectively captured by refining a single layer rather than modifying the entire model. Identifying the optimal layer could further enhance performance. On the other hand, training multiple layers at once, either from the early stage (layers 1–6) or the later stage (layers 7–12), leads to a significant drop in performance. This suggests that updating too many layers disrupts the task-agnostic knowledge embedded in the pre-trained model, reducing its ability to generalize across tasks.
@@
Figure 4: Impact of training different task-specific layer. We compare the average performance of merged models when different task-specific layers are trainable, while other layers update only their merging coefficients. Training a single layer achieves accuracy similar to tuning the classifier, whereas modifying multiple layers leads to performance degradation.
Sparsity analysis of merging coefficients. Inspired by the literature [75, 72, 11], which highlights how redundant parameters degrade performance due to conflicts among task-specific parameters, we investigate this phenomenon in our method. We analyze the learned merging coefficients to observe their impact, specifically exploring sparsity after training them alone and jointly with the classifier. We find 37.2% versus 55.9% of the coefficients are concentrated near zero (i.e., smaller than 1e-5) with the deviated average accuracy of 83.4% versus 89.6% for each configuration. This suggests that training a task-specific layer, like the classifier, along with merging coefficients, reduces task interference by pruning unnecessary parameters, leading to improved accuracy. For visualizations of merging coefficients, please refer to Appendix D.4.
@@
Initialization robustness. Our method benefits from robust merging coefficient initialization. To highlight this, we present training curves under different initialization settings and compare them with those of other model merging methods. As shown in Figure 5, when only the classifier is trained (i.e., fixed shared encoder), the performance is sensitive to initialization; However, training both the classifier and coefficients together improves robustness. Figure 6 shows that when training a task-specific adapter only, the Surgery model is sensitive to initialization. Unlikely, our method jointly optimizes the merging coefficient and task-specific layer, making it robust to initialization sensitivity.
@@ 6 Conclusion
In this paper, we have introduced Model Tinting , an efficient multi-task model merging approach that enhances task-specific knowledge with minimal overhead. By adaptively learning task-specific layers while refining the shared model, our method balances generalization and specialization without additional task-specific modules. This distinguishes our method from conventional model merging methods, which often introduce unnecessary complexity or struggle to retain fine-grained task information. To further improve robustness, Model Tinting++ incorporates a confidence-based filtering mechanism that selectively removes unreliable supervision signals. Extensive experiments across vision, NLP, and dense prediction tasks demonstrate that our approach consistently achieves state-of-the-art performance, particularly excelling in dense prediction and large-scale multi-task benchmarks. Additionally, our analysis reveals key insights: 1) Inner layers complement each other in capturing task-specific features; 2) Training task-specific layers induces sparsity in merging coefficients, improving task separation; 3) Our joint optimization strategy is robust to initialization, ensuring stable results across tasks. We believe this work provides a strong foundation for future research in scalable and adaptive model merging.
@@
  • Appendix C: Additional results for pilot study, NLP, and vision tasks;
@@ Appendix A Overview of Model Tinting
We give a brief overview of our proposed method Model Tinting , a test-time method that merges multiple fine-tuned models using a single trainable layer per task Figure A provides an illustrative overview of our method. At the top, the figure illustrates the merging process, where multiple fine-tuned models ($\phi_{1},\phi_{2},\ldots,\phi_{K}$) are combined into a shared representation ($\phi_{\text{MTL}}$) using adaptive merging coefficients. During this process, a trainable task-specific layer ($\phi_{\text{tint}}^{l}$) is introduced at a chosen transformer layer $l$, capturing task-specific information from the task-agnostic representations of the merged encoder. We argued that any layer can be picked for task adjustments before merging, and this is further supported by the graph in Figure A. The results show that applying Model Tinting, regardless of the chosen layer, leads to significant performance improvements over the baseline. This highlights the robustness and effectiveness of task-specific adjustments in enhancing task-agnostic representations, with more details provided in Appendix D.1. We also proposed Model Tinting++ , which applies a strategy to filter noisy predictions from individual models.
@@ B.1 Training Details
Main experiments. We initialize the layer-wise merging coefficients for all layers to 0.3 by default, following the values used in the previous method [74]. However, for the 14 and 20 vision tasks, the coefficients are initially set to 0.1, as using 0.3 incurred significantly lower Task Arithmetic [23] performance for these tasks, presumably due to the increased number of tasks. We use the Adam [29] optimizer with momentum parameters (0.9, 0.999) to update the coefficients and the classifier. For vision tasks, the learning rate is set to 0.01 when training the classifier and 0.001 when training only the merging coefficients. For dense prediction tasks, the learning rate of 0.0001 is used. For NLP tasks, the learning rate of 0.0005 is used, regardless of whether the classifier, the merging coefficients, or both are being updated. Consistent with prior works [74, 58, 73], vision tasks are trained with 16,000 test samples per task over 500 iterations, using a batch size of 32. Dense prediction tasks use 400 samples per task over 25 iterations with a batch size of 16. NLP tasks are trained with 32,000 samples per task over 1,000 iterations with a batch size of 32. For each task, the loss is computed, and the model parameters are updated immediately following the forward pass for each batch. To enable sequential processing, the order of tasks is randomized.
To address this, we experiment with a sequential update strategy, where updates are applied immediately after each task’s forward pass instead of the typical one. Interestingly, this sequential update approach gives a positive byproduct: performance improvements over the traditional approach. While the original AdaMerging achieves an average accuracy of 80.1%, our sequential update strategy improves it to 81.6% for the 8 vision tasks using ViT-B/32. Similarly, our method also enjoys a gain when using sequential updates (89.6% compared to 89.4%). We argue that this improvement may result from mitigating catastrophic forgetting in continual learning [30] by training tasks sequentially rather than simultaneously.
Pilot study. The aforementioned training details are applied almost identically to the experiments in the pilot study of our main paper. As stated in the main paper, only the classifier was chosen (as a straightforward and practical option) to train on the given Task Arithmetic (or AdaMerging [74]) merged weights for the 8 vision tasks using ViT-B/32. A difference is that, instead of using the test dataset, the training dataset is used in a supervised manner to produce the results closer to the upper bound, in which the difference stands out more. We train the models for only one epoch.